If 3.03 m3m3 of a gas initially at STP is placed under a pressure of 3.49 atm , the temperature of the gas rises to 36.3 ∘C. What is the volume?
Answer #1

initial volume = V1 = 3.03 m^3

initial pressure = P1 = 1 atm

initial temperature = T1 = 273 k

final temperature = T2 = 36.3 + 273 =309.3 K

final pressure = P2 = 3.49 atm

                     P1 * V1 / T1 = P2 * V2 / T2

                 V2 = P1* V1* T2 / (P2* T1)

                  V2 = 1 * 3.03 * 309.3 / ( 3.49* 273)

                  V2 = 0.98 m^3